# Optimizing for ROC Curves on Class-Imbalanced Data by Training over a Family of Loss Functions

### v1

@@ 6.5 Choice of $\lambda$
In this section, we compare different options for which hyperparameters in VS loss to use as $\boldsymbol{\lambda}$ in LCT. We consider setting $\boldsymbol{\lambda}$ to each of the three hyperparameters (using constant values for the other two). We also consider setting $\boldsymbol{\lambda}=(\Omega,\gamma,\tau)$. Figure 8 compares these options and shows that $\boldsymbol{\lambda}=\tau$ achieves a good compromise between improving the AUC of the models and reducing the variance of results. Appendix F compares $\boldsymbol{\lambda}=\tau$ and $\boldsymbol{\lambda}=(\Omega,\gamma,\tau)$ on the Melanoma dataset and finds similar results. Furthermore, this choice of $\lambda$ follows our insights from Section 4.
Figure 8: LCT with different $\boldsymbol{\lambda}$. Each boxplot contains results for 27 models with different hyperparameter values. All models were trained on CIFAR10 cat/dog with $\beta=100$.
@@ 7 Conclusion
In conclusion, we propose using Loss Conditional Training (LCT) to train binary classification models under severe class imbalance. We find that this consistently both improves the ROC curves and reduces the model’s sensitivity to hyperparameter choices. This improvement comes from the fact that training over a family of loss functions is a proxy for optimizing along different TPR and FPR tradeoffs. Areas of future work include studying how to adapt this method to work on multi-class classification problems under imbalance and regression tasks.
@@ Appendix A Binary classification metrics
[…]
@@ A.1 True Positive Rate (TPR) = Minority-class accuracy = Recall
[…]
@@ A.2 False Positive Rate (FPR) = 1 - Majority-class accuracy = 1 - TNR
[…]
@@ A.3 Precision
[…]
@@ A.4 Overall Accuracy
[…]
@@ A.5 $F_{1}$ and $F_{\beta}$
[…]
@@ A.6 G-mean
[…]
@@ A.7 ROC curves and AUC
Of course, the number of true positives and true negatives are a trade-off and any method can be modified to give a different combination of these metrics. Specifically, the decision threshold can be modified to give any particular recall. Receiver Operating Characteristic (ROC) curves take this in consideration and show the trade-off of true positive rates and false positive rates over all possible decision thresholds. The area under the ROC curve (ROC AUC) is calculated using the trapezoid rule for integration over a sample of values and is a commonly used metric under class imbalance (Buda et al., 2018).
Appendix B Detailed related work
Training on imbalanced datasets with algorithms designed to work on balanced datasets can be problematic because the gradients and losses are biased towards the common classes, so the rare classes will not be learned well. Current methods to mitigate the effects of training under imbalance include methods that modify the loss functions, re-sample and augment training samples, and improve the module via two-stage learning, ensembles, or representation learning (Zhang et al., 2023).
B.1 Specialized loss functions
To balance the gradients from all classes, several papers adaptively change a sample’s weight in the loss based on features such as the sample’s confidence score, class frequency, and influence on model weights (Zhang et al., 2021; Fernando & Tsokos, 2021; Park et al., 2021; Wang et al., 2021a; Li et al., 2021a). Other work addresses the difference in the norms of features associated with frequent (head) and rare (tail) classes and proposes to balance this by utilizing a feature-based loss (Li et al., 2022b) or weight-decay and gradient clipping (Alshammari et al., 2022).
Some work has focused on enforcing larger margins on the tail classes using additive factors on the logits in Cross-entropy loss (Cao et al., 2019; Menon et al., 2021; Li et al., 2022c). Other work proposed adding multiplicative factors to the logits to adjust for the difference in the magnitude of minority-class logits at training and test time or minimize a margin-based generalization bound (Ye et al., 2020; Kang et al., 2021). Kini et al. (2021) show that multiplicative factors are essential for the terminal phase of training, but that these have negative effects early during training, so additive factors are necessary to speed up convergence. They propose Vector Scaling (VS) loss as a general loss function that includes both additive and multiplicative factors on the logits. Behnia et al. (2023) study the implicit geometry of classifiers trained on a special case of VS loss.
Beyond classification, many papers have focused on imbalance in instance segmentation and object detection applications (Tan et al., 2021; Wang et al., 2021c; Feng et al., 2021; Li et al., 2020).  Ren et al. (2022) also propose a balanced Mean Square Error (MSE) loss for regression problems, such as age and depth estimation.
B.2 Data-level methods
Another way to balance the gradients of classes during training is resampling. This could be done by sampling the minority class more often (random over-sampling) or sampling the majority class less often (random under-sampling) (Johnson & Khoshgoftaar, 2019; Japkowicz, 2000; Liu et al., 2009). Jiang et al. (2023); Hou et al. (2023) use clustering to drive resampling; specifically, they cluster head classes into multiple clusters and then resample across clusters instead of classes. Meta-learning has been used to estimate the optimal sampling rates of different classes (Ren et al., 2020), while Wang et al. (2019) dynamically adapts both the loss function and sampling procedure throughout training.
Additionally, data augmentation can be used alongside oversampling to increase the size of the minority class samples and enable model generalization (Zhong et al., 2021; Zang et al., 2021; Li et al., 2021b; Du et al., 2023). With tabular data, small perturbations of random noise can be added to generate new examples. Images lend themselves to more high-level augmentations. Methods that copy and paste patches of images, such as CutMix (Yun et al., 2019) and PuzzleMix (Kim et al., 2020) have been used to improve classification or instance segmentation (Ghiasi et al., 2020) performance.
SMOTE (Synthetic Minority Over-sampling Technique) creates synthetic examples by interpolating between samples in the same class of the training set (Chawla et al., 2011). The interpolation is done in feature space instead of data space. The mixup method also generates synthetic examples; however, unlike SMOTE, it interpolates between samples of different classes (Zhang et al., 2018). These synthetic examples are given a soft label that corresponds to the proportion of input from each class. StyleMix adapts mixup to separately manipulate an image’s content and style (Hong et al., 2021). ReMix advances mixup to optimize for situations with class imbalance by combining mixup and resampling (Bellinger et al., 2020). Li et al. (2021b) creates augmented data samples by translating deep features along semantically meaningful directions. Zada et al. (2022) adds pure noise images to the training set and introduces a new type of distribution-aware batch normalization layer to facilitate training with them.
B.3 Module improvements
Some work has shown that overly emphasizing the minority class can disrupt the original data distribution and cause overfitting (Zhou et al., 2020; Du et al., 2023) or reduce the accuracy of the majority class (Zhu et al., 2022). To mitigate this issue, contrastive learning (Du et al., 2023; Zhu et al., 2022; Wang et al., 2021b) has been used to learn a robust feature representation by enforcing consistency between two differently augmented versions of the data. Du et al. (2023) devise a single-stage approach, by combining the ideas of contrastive representations with soft label class re-weighting to achieve state-of-the-art performance on several benchmarks.
Multi-expert methods have also been explored (Zhang et al., 2022; Wang et al., 2020b; Li et al., 2022a).  Li et al. (2022a) collaboratively learn multiple experts, by learning individual experts and transferring knowledge among them, in a nested way. Methods have also used two branches or heads with one classification branch along with another branch, either to re-balance (Guo & Wang, 2021; Zhou et al., 2020) or calibrate (Wang et al., 2020a) the model. Tang et al. (2022) adopt a generalized approach to learn attribute-invariant features that first discovers sets of samples with diverse intra-class distributions from the low confidence predictions, and then learns invariant features across them. Dong et al. (2022) uses language prompts to improve both general and group-specific features.
Appendix C Number of samples
@@
Appendix D Effect of evaluation $\lambda$
Figure 9: Effect of $\boldsymbol{\lambda}$ value at inference time. Results are shown for one LCT model that has $\boldsymbol{\lambda}=\tau$. Model was trained using the protocol described in Section 6 with $\boldsymbol{\lambda}=\tau$ drawn from a linear distribution on $[0,3]$ with $h_{b}=0$. Model is evaluated at four $\boldsymbol{\lambda}=\tau$ values: $0.0,1.0,2.0$, and $3.0$. For each evaluation $\boldsymbol{\lambda}$, we plot a) the model’s TPR/FPR when $t=0.5$ (circle) and b) the ROC curve of the model (curve). The evaluation $\boldsymbol{\lambda}$ affects the TPR/FPR for a constant $t$ (e.g., $t=0.5$); however, the ROC curves of models evaluated on different $\boldsymbol{\lambda}$s are almost identical.
Appendix E Linear probability density function
@@
Appendix F Using $\lambda=(\Omega,\gamma,\tau)$ on the SSIM Melanoma dataset
Figure 11: Comparison of baseline, LCT with $\boldsymbol{\lambda}=\tau$, and LCT with $\boldsymbol{\lambda}=(\Omega,\gamma,\tau$ on the Melanoma dataset with $\beta=100.$ The two LCT methods are comparable with $\boldsymbol{\lambda}=\tau$ having a slight edge over $\boldsymbol{\lambda}=(\Omega,\gamma,\tau$ in terms of the best model at high TPRs.
Appendix G Simplifying VS Loss
@@
Appendix H Partial derivatives of VS-Loss with respect to $\Omega,\tau,\gamma$
@@
Appendix I Full derivation for break-even points
@@ I.1 Numerical Example
Appendix J Contour plots of effect of $\Omega,\gamma,\tau$ on VS loss
@@
J.1 Regular cross-entropy
@@
J.2 Varying $\Omega$ (Weighted cross-entropy loss)
Next consider VS loss with arbitrary weights $\Omega$, but fixed $\gamma=0,\tau=0$. Note that this is equivalent to weighted cross-entropy with class 1 having weight $\Omega$ and class 0 having weight $1-\Omega$. We include contour plots of this case in the top row of Figure 12 and simplify the equation of equal losses below. Unfortunately there is no analytic solution in this case, but we provide more insights about this case in Subsection J.5.
@@
J.5 $\Omega$ as a function of softmax scores
@@
$p_{1}$ $=\frac{e^{\Delta_{1}z_{1}+\iota_{1}}}{e^{\Delta_{0}z_{0}+\iota_{0}}+e^{\Delta_{1}z_{1}+\iota_{1}}}$ (52)
@@
Because 0-1 loss is not differentiable, models are instead optimized for a proxy function like cross-entropy. In traditional cross-entropy (i.e., $\Omega=0.5$), the loss function is symmetric such that $\ell(0,p_{1})=\ell(1,-p_{1})$ which results in the DNN being optimal for a threshold of 0.5 of $p_{1}$. Figure 13 shows how this threshold varies for different values of $\Omega$. As $\Omega$ increases, the line $y=1$ becomes steeper while $y=0$ becomes less steep and the value of $p_{1}$ where the losses intersect increases. Note that the point of intersection is
@@
Additionally, note that changing $\gamma$ or $\tau$ only affects the calculation of $p_{1}$, but does not change the point $p_{1}$ where the losses intersect.

### v2

@@ 6.5 Choice of $\lambda$
In this section, we compare different options for which hyperparameters in VS loss to use as $\bm{\lambda}$ in LCT. We consider setting $\bm{\lambda}$ to each of the three hyperparameters (using constant values for the other two). We also consider setting $\bm{\lambda}=(\Omega,\gamma,\tau)$. Figure 8 compares these options and shows that $\bm{\lambda}=\tau$ achieves a good compromise between improving the AUC of the models and reducing the variance of results. Furthermore, this choice of $\lambda$ follows our insights from Section 4.
@@ 7 Conclusion
In conclusion, we analyze the theoretically well-studied Vector Scaling (VS) loss, and identify that, in practice, its performance exhibits significant variability for binary class imbalance problems with severe imbalance. We mitigate this problem by training over a family of loss functions. We find that this consistently both improves the ROC curves and reduces the model’s sensitivity to hyperparameter choices. This improvement comes from the fact that training over a family of loss functions is a proxy for optimizing along different TPR and FPR tradeoffs. Areas of future work include studying how to adapt this method to work on multi-class classification problems under imbalance and regression tasks.
@@ Appendix A Binary classification metrics
Actually Positive True Positives (TP) False Negatives (FN)
Actually Negative False Positives (FP) True Negatives (TN)
@@ A.1 True Positive Rate (TPR) = Minority-class accuracy = Recall
Actually Positive True Positives (TP) False Negatives (FN)
@@ A.2 False Positive Rate (FPR) = 1 - Majority-class accuracy = 1 - TNR
Actually Negative False Positives (FP) True Negatives (TN)
@@ A.3 Precision
Actually Positive True Positives (TP) False Negatives (FN)
Actually Negative False Positives (FP) True Negatives (TN)
@@ A.4 Overall Accuracy
Actually Positive True Positives (TP) False Negatives (FN)
Actually Negative False Positives (FP) True Negatives (TN)
@@ A.5 $F_{1}$ and $F_{\beta}$
Actually Positive True Positives (TP) False Negatives (FN)
Actually Negative False Positives (FP) True Negatives (TN)
@@ A.6 G-mean
Actually Positive True Positives (TP) False Negatives (FN)
Actually Negative False Positives (FP) True Negatives (TN)
@@ A.7 ROC curves and AUROC
Of course, the number of true positives and true negatives are a trade-off and any method can be modified to give a different combination of these metrics. Specifically, the decision threshold can be modified to give any particular recall. Receiver Operating Characteristic (ROC) curves take this in consideration and show the trade-off of true positive rates and false positive rates over all possible decision thresholds. The area under the ROC curve (AUROC) is calculated using the trapezoid rule for integration over a sample of values and is a commonly used metric Buda et al. (2018).
A.8 Precision-Recall Curves and AUPRC
Similarly, the precision-recall curves show the tradeoff of the recall (on the x-axis) and the precision (on the y-axis) over all possible thresholds. The area under the Precision-Recall Curve (AUPRC) is also calculated using the trapezoid rule.
Appendix B Number of samples
@@
Appendix C Effect of evaluation $\lambda$
Figure 9: Effect of $\bm{\lambda}$ value at inference time. Results are shown for one LCT model that has $\bm{\lambda}=\tau$. Model was trained using the protocol described in Section 6 with $\bm{\lambda}=\tau$ drawn from a linear distribution on $[0,3]$ with $h_{b}=0$. Model is evaluated at four $\bm{\lambda}=\tau$ values: $0.0,1.0,2.0$, and $3.0$. For each evaluation $\bm{\lambda}$, we plot a) the model’s TPR/FPR when $t=0.5$ (circle) and b) the ROC curve of the model (curve). The evaluation $\bm{\lambda}$ affects the TPR/FPR for a constant $t$ (e.g., $t=0.5$); however, the ROC curves of models evaluated on different $\bm{\lambda}$s are almost identical.
Appendix D Linear probability density function
@@
Appendix E Simplifying VS Loss
@@
Appendix F Partial derivatives of VS-Loss with respect to $\Omega,\tau,\gamma$
@@
Appendix G Full derivation for break-even points
@@ G.1 Numerical Example
Appendix H Contour plots of effect of $\Omega,\gamma,\tau$ on VS loss
@@
H.1 Regular cross-entropy
@@
H.2 Varying $\Omega$ (Weighted cross-entropy loss)
Next consider VS loss with arbitrary weights $\Omega$, but fixed $\gamma=0,\tau=0$. Note that this is equivalent to weighted cross-entropy with class 1 having weight $\Omega$ and class 0 having weight $1-\Omega$. We include contour plots of this case in the top row of Figure 11 and simplify the equation of equal losses below. Unfortunately there is no analytic solution in this case, but we provide more insights about this case in Subsection H.5.
@@ H.4 Varying $\tau$
H.5 $\Omega$ as a function of softmax scores
@@
$p_{1}$ $=\frac{e^{z_{1}}}{e^{z_{0}}+e^{z_{1}}}$ (52)
@@
Because 0-1 loss is not differentiable, models are instead optimized for a proxy function like cross-entropy. In traditional cross-entropy (i.e., $\Omega=0.5$), the loss function is symmetric such that $\ell(0,p_{1})=\ell(1,-p_{1})$. Figure 12 shows how this threshold varies for different values of $\Omega$. As $\Omega$ increases, the line $y=1$ becomes steeper while $y=0$ becomes less steep and the value of $p_{1}$ where the losses intersect increases. Note that the point of intersection is
@@
Additionally, note that changing $\gamma$ or $\tau$ does not change the point $p_{1}$ where the losses intersect.